OpenStudy (anonymous):

f(z)=tanh x+i tanh y what is the domain for the function f(z) is analytic and where it is not analytic?

OpenStudy (anonymous):

sinh x = (ex-ex)/2 cosh x = (ex+ex)/2 tanh x = sinhx/coshx now check limits for denomitor not equal zero and that will be its domain and go for mind that wat will be range limit it will be cleared